OpenStudy (anonymous):

Solve x^2 - 8x + 3 = 0 by completing the square. Which equation is used in the process? a. (x- 4)^2= 0 b.(x- 4)^2= 13 c. (x- 4)^2= 19

jhonyy9 (jhonyy9):

add to both sides 13

OpenStudy (ajprincess):

x^2 - 8x + 3 = 0 x^2-8x+16+3-16=0 x^2-8x+16-13=0 x^2-4x-4x+16=13 x(x-4)-4(x-4)=13 (x-4)(x-4)=13 (x-4)^2=13 Is that clear?

OpenStudy (anonymous):

answer is "c" did u get it?

jhonyy9 (jhonyy9):

b is right

OpenStudy (anonymous):

s "b"